OpenStudy (anonymous):

graph |x-3| is greater than and equal to 5 on a number line

zepdrix (zepdrix):

\[\Large\rm |x+3|\ge 5\]\[\Large\rm |x+3|=\pm(x+3)\]Let's plug this in,\[\Large\rm \pm(x+3)\ge5\]We'll have two equations,\[\Large\rm +(x+3)\ge5,\qquad\qquad -(x+3)\ge5\]

zepdrix (zepdrix):

Let's solve for x in each case.

zepdrix (zepdrix):

Ahh it's minus 3, darn it! Sorry bout that.\[\Large\rm +(x-3)\ge5,\qquad\qquad -(x-3)\ge5\]

OpenStudy (anonymous):

so x will equal\[x \ge8 and x \ge2\]

zepdrix (zepdrix):

So your solution to the first inequality looks good! Let's work on the second one... see what went wrong.

OpenStudy (anonymous):

whoops i forgot to put the x negative

zepdrix (zepdrix):

\[\Large\rm -(x-3)\ge5\]Multiply both sides by -1,\[\Large\rm x-3\le-5\]Remember that your inequality `flips` when you toss a negative across, yes?

OpenStudy (anonymous):

so now it's \[x \ge8 and x \le-8\]

zepdrix (zepdrix):

So remember, we're `adding` 3, not subtracting.

zepdrix (zepdrix):

The sign didn't change on the 3 since we multiplied the -1 to the other side.

OpenStudy (anonymous):

oh my goodness i feel dumb now haha! i swear I'm actually good at math!

zepdrix (zepdrix):

hehe

OpenStudy (anonymous):

so then it will -2

zepdrix (zepdrix):

yaayyy good job \c:/